OpenStudy (anonymous):

Calculus: Is this correct? All edges of a cube are expanding at a rate of 6 cm/sec. How fast is the volume changing when each edge is 2 cm? V=x^3 V'=3x^2(x') V(2)=3(2cm)^2(6cm/sec) V(2)=72cm^3/sec I am not sure I did this correctly, but I think I did...

OpenStudy (kenljw):

V = (x(0)+vt)^3 if x(0) =0 then V=(vt)^3=2^3 V'=3(vt)^2 v V'=3(2)^2 6=72

OpenStudy (anonymous):

I don't have a great understanding of this. Can you explain a little? This looks like you are refuting my answer?

OpenStudy (anonymous):

I think I got it... There should be a multiplication between ^2 and your 6?

OpenStudy (kenljw):

Eq 1 is volume with arbitrary initial side length Eq 2 say's initial side length 0 Eq 3 is resulting volume at side length 2 Eq 4 takes the derivative of volume, v is constant therefore v' is 0 Eq 5 place result of Eq 3 and solved

OpenStudy (radar):

I think you did too.

OpenStudy (anonymous):

Thanks guys!